# Privacy-Preserving Video Anomaly Detection: A Survey

### v1

@@ III-A Encrypted Videos-based P2VAD
TABLE II: Encrypted Video Coding Techniques for P2VAD.
@@
Encrypted and compressed videos obscure sensitive appearance information recognizable by humans and filter out low-frequency data, making them inherently suited for detecting anomalous behaviors tied to temporal cues. Additionally, compressed videos require much less storage and bandwidth than raw RGB sequences, reducing transmission costs and enabling real-time VAD. The available encrypted video coding techniques are summarized in Table II. For instance, Biswas and Babu [89] found that motion vector amplitudes in the H.264/AVC compressed domain could be effectively used to simulate usual movement patterns and detect anomalies. This method not only addresses privacy concerns but also achieves processing speeds of over 150 frames per second (fps) on the UMN [90] and UCSD Peds [91] datasets, with up to 200$\times$ acceleration. They also introduced a motion pyramid-based hierarchical approach to capture temporal dynamics more efficiently, achieving detection accuracy comparable to state-of-the-art pixel-domain algorithms. In a follow-up study [70], they enhanced the feature set by including motion vector orientation, further improving anomaly detection and achieving 90$\times$ and 250$\times$ speedups on two datasets, respectively. Another method, based on the Histogram of Oriented Motion Vectors (HOMV) within a sparse representation framework [92], utilizes online dictionary learning to represent the sparse behavior of HOMVs. It detects anomalies by analyzing the likelihood of sparse coefficients occurring at specific locations in the H.264 compressed video.
@@ IV Desensitized Intermediate Modalities P2VAD
Newer techniques, such as multi-task learning [100] and causality learning [101, 102, 103], have been explored in RGB-based VAD models, allowing for more specialized categories. This progress is expected to guide future skeleton-based VAD research. In contrast, optical flow removes scene appearance information and is less information-dense than skeletal data. Therefore, it is often used as complementary semantics to RGB sequences and is not typically applied as a standalone VAD approach. Studies have shown that such intermediate modalities, which focus on temporal information, not only mitigate privacy concerns but also avoid enrollment sensitivities similar to those in RGB-based methods. In real-world scenarios, factors such as scene changes, camera angles, and weather conditions cause significant pixel-level variations in color images, leading to high false positive rates in unsupervised VAD methods that rely solely on normal video modeling. Scene-robust VAD has thus become a long-standing challenge. The Desensitized Intermediate Modalities (DIM) approach, which focuses on motion information while ignoring background scenes, allows VAD to concentrate on learning scene-robust video normality.
@@ V Edge-Cloud Intelligence Empowered P2VAD
With the growing adoption of video IoT systems in smart cities and the expansion of online video platforms, data transmission and processing in VAD systems have transitioned from centralized local devices to distributed hybrid architectures that combine edge and cloud computing. Edge-cloud intelligence-enabled P2VAD research focuses on leveraging privacy computing and distributed privacy-preserving techniques to enhance the security of VAD systems across end devices, edge nodes, and the cloud. Unlike the approaches presented in Sections III and IV, which focus on privacy risks during data acquisition and preprocessing by removing identifiable appearance information, Edge-Cloud Intelligence (ECI) empowerd VAD emphasizes system-level data security and privacy preservation [127]. Although localized VAD models may still use RGB sequences as inputs, these computational processes typically occur on user-trusted devices, such as smartphones or PCs. Instead of collecting raw data, data centers and cloud servers process encrypted information or intermediate features, ensuring that even in the event of a system breach, sensitive information remains secure. Furthermore, the distributed processing of surveillance videos from different scenes prevents distribution bias caused by simple data aggregation, allowing cloud models to perform cross-scene anomaly detection.
@@
VAD systems for large-scale applications handle massive amounts of video data, which typically need to be transmitted and processed across various edge devices, servers, and data centers. Without effective privacy protection mechanisms, the exchange and integration of information in edge-cloud systems pose significant risks of data leakage and misuse. Therefore, system-level privacy protection not only safeguards user data but also strengthens societal trust in such systems. This chapter explores three major directions: (1) Privacy Computing-enhanced VAD, which reduces the risk of data leakage through privacy computing techniques; (2) Edge-Cloud Collaboration VAD, which focuses on data security and performance optimization during collaboration between end devices and the cloud; and (3) Federated VAD, which explores privacy-secure VAD model training via distributed learning without sharing raw data. These approaches complement one another and collectively improve system-level privacy security.
@@ V-B Edge-Cloud Collaboration VAD
Edge-Cloud Collaboration VAD is a distributed architecture that strategically divides computing tasks between edge devices and the cloud, aiming to maximize the use of computing resources and alleviate users’ privacy and security concerns about data centers [130]. In this framework, video data captured from real-world scenes can be initially processed on edge devices close to the data source, and only essential features or privacy-protected data are sent to the cloud for further analysis. Since only user-trusted local devices access raw data containing identifiable information, the end-edge-cloud data exchange process and global model training can avoid the risk of data theft and misuse. Existing research in this area is typically categorized into three dimensions: (1) Edge reasoning: smartphones and PCs deployed on the end-side utilize local computing power to preprocess video streams and detect simple anomalies in real-time. Lightweight deep VAD models are often used, reducing data transmission and minimizing privacy exposure; (2) Cloud aggregation: the cloud collects intermediate outputs from the edge, such as hidden features or decision vectors devoid of identifiable information, to train cross-scene models for detecting complex, long-duration anomalies; and (3) Secure data transmission: to prevent the decoding of intermediate features into identity-identifiable information, data is encrypted during transmission across devices, ensuring that sensitive information is not leaked.
@@ VI-A1 Generalized VAD Datasets
TABLE III: Statistical information of the generalized VAD datasets.
@@
The generalized datasets contain RGB video sequences captured from real-world scenes and are commonly used for traditional deep VAD model training and testing. The details are presented in Table III. In P2VAD models, these datasets are often encoded or transformed, with morphological abstraction, optical flow extraction, or skeleton detection applied.
@@ VI-B3 Privacy-Preservation Oriented Metrics
  • Degree of De-identification: This measures the effectiveness of techniques like visual desensitization and feature abstraction in removing identifiable information. This can be indirectly assessed through image quality metrics such as Structural Similarity Index (SSIM) or Peak Signal-to-Noise Ratio (PSNR). A higher degree of de-identification corresponds to lower similarity between the processed and original content.
@@ VII Discussion
TABLE IV: Comparison of P2VAD Pathways.
@@
P2VAD with Non-Identifiable Elements In the post-data acquisition step, encryption compression, appearance abstraction, or NVLCs are used to remove personally identifiable information from the RGB sequence. Eliminates privacy concerns and enhances transparency at the data source. Encryption compression and appearance abstraction can work on existing videos. Encrypted video reduces redundant background and increases inference speed by over 50 times. NVLCs can function in extreme conditions, such as darkness and high-speed environments. Encryption retains only high-frequency components, limiting detection performance of appearance-motion joint anomalies. Performance of appearance abstraction depends on additional methods (e.g., object detection and instance segmentation). Non-visible light videos require specialized sensing equipment, increasing hardware and data preparation costs. The ideas of RGB-based VAD methods are often incompatible with such data. Efficient video compression can enable the real-time P2VAD system. Lightweight object detection and zero-shot segmentation models (e.g., SAM) can handle complex scenes. Sensing and representation technologies with event cameras and infrared will further P2VAD development. AI technologies (e.g., adversarial samples) can enhance data security and the robustness of P2VAD models.
@@
Sections III-V have presented the implementation and development of various categories of P2VAD approaches that are motivated by different research objectives and underlying assumptions. While existing works have made notable progress in data collection, model learning, and system deployment, significantly advancing privacy-preserving techniques in video anomaly detection, P2VAD, as an emerging interdisciplinary research topic, still faces unresolved challenges in real-world applications. Furthermore, the proliferation of intelligent video surveillance systems and the widespread popularity of short video applications (e.g., TikTok, Kuaishou, and Youtube) among Generation Z provide a broad market for P2VAD. Meanwhile, advancements in AI security and IoT technologies are expected to deeply integrate with VAD research, significantly driving the development of P2VAD. The brief comparison of the strengths, weaknesses, and future directions of various P2VAD Pathways are outlined in Table IV. Moreover, from the perspective of privacy preservation and the research of trustworthy P2VAD systems for large-scale video IoT, this section discusses the challenges and opportunities of P2VAD, analyzing the limitations of existing approaches and trends for future exploration.
@@ VII-A2 Computational Overhead in Heterogeneous Systems
Urban-scale VAD systems often rely on edge intelligence techniques (e.g., mobile computing, edge-cloud collaboration) to optimize the allocation of computational resources and improve overall performance. However, addressing privacy preservation in such large-scale heterogeneous systems—comprising numerous devices such as surveillance cameras, smartphones, local gateways, data centers, and servers—not only requires efficient privacy-preserving computing strategies but also faces significant increases in computational cost. Edge computing typically requires sensitive raw RGB sequences to be processed locally on resource-constrained yet trusted client devices. Additionally, intermediate features often need to be encrypted before transmission to servers or global models to prevent data misuse during transmission. This increases the computational burden on local devices and affects the system’s real-time detection performance. Deploying privacy-compliant VAD systems in scenarios requiring fast response and large-scale application remains a bottleneck, necessitating the development of specialized privacy-preserving computing techniques for video data and efficient computation offloading algorithms for heterogeneous video IoT devices.
@@ VII-A4 Common Challenges from Emerging AI Technologies
Large models (e.g., Generative Pre-Training [149] and Contrastive Language-Image Pretraining [150]) have been introduced to enhance VAD’s ability to understand and describe anomalies [151, 152, 153]. However, their human-like interaction capabilities [154] have further exacerbated public concerns about privacy in VAD. For example, when capturing supplementary semantics in RGB video sequences to improve VAD performance, large models tend to fully describe all visual information, including facial details, clothing appearance, gender, and ethnicity, without considering privacy implications [155]. While spatial appearance and temporal motion cues are essential for VAD tasks, the presence of sensitive identity information is typically irrelevant and exacerbates public distrust. Furthermore, Embodied Artificial Intelligence (EAI) [156] and Explainable Machine Learning (XML) introduce new scenarios and research perspectives for P2VAD. EAI combines robots and perception devices capable of real-time environmental sensing and autonomous responses to emergencies, yet these systems must process large amounts of sensor data (e.g., RGB cameras, infrared cameras, and event cameras), making privacy concerns prominent. XML improves the transparency of P2VAD systems by allowing users to understand the model’s detection process and decision logic. However, it may expose sensitive information, raising questions about balancing explainability and privacy preservation.

### v2

@@ III-A Encrypted Videos-based P2VAD
TABLE III: Encrypted Video Coding Techniques for P2VAD.
@@
Encrypted and compressed videos obscure sensitive appearance information recognizable by humans and filter out low-frequency data, making them inherently suited for detecting anomalous behaviors tied to temporal cues. Additionally, compressed videos require much less storage and bandwidth than raw RGB sequences, reducing transmission costs and enabling real-time VAD. The available encrypted video coding techniques are summarized in Table III. For instance, Biswas and Babu [97] found that motion vector amplitudes in the H.264/AVC compressed domain could be effectively used to simulate usual movement patterns and detect anomalies. This method not only addresses privacy concerns but also achieves processing speeds of over 150 frames per second (fps) on the UMN [103] and UCSD Peds [104] datasets, with up to 200$\times$ acceleration. They also introduced a motion pyramid-based hierarchical approach to capture temporal dynamics more efficiently, achieving detection accuracy comparable to state-of-the-art pixel-domain algorithms. In a follow-up study [70], they enhanced the feature set by including motion vector orientation, further improving anomaly detection and achieving 90$\times$ and 250$\times$ speedups on two datasets, respectively. Another method, based on the Histogram of Oriented Motion Vectors (HOMV) within a sparse representation framework [98], utilizes online dictionary learning to represent the sparse behavior of HOMVs. It detects anomalies by analyzing the likelihood of sparse coefficients occurring at specific locations in the H.264 compressed video.
@@ IV Desensitized Intermediate Modalities P2VAD
TABLE IV: Technical Comparison of DIM-based P2VAD Methods
Method Year Category Core Contributions Feature Type Learning Framework Network Architecture
Gatt et al. [108] 2019 Skeleton-Agent LSTM+CNN autoencoder for normal event representation Body keypoints Reconstruction LSTM+CNN
Markovitz et al. [109] 2020 Skeleton-Clustering Graph Embedded Pose Clustering with GCAE Human pose maps Deep clustering GCAE
Rodrigues et al. [110] 2020 Skeleton-Agent Multi-timescale model for pose trajectory prediction Pose trajectories Prediction Multi-timescale model
Temuroglu et al. [111] 2020 Skeleton-Agent Occlusion-Aware Skeleton Trajectory Representation Skeleton trajectories Reconstruction Autoencoder
Cui et al. [112] 2021 Skeleton-Clustering Prototype generation module with SST-GCN Skeleton features Deep clustering SST-GCN
HSTGCNN [113] 2021 Skeleton-Agent Hierarchical high/low-level graph representations Graph representations Prediction ST-GCN
Liu et al. [114] 2021 Skeleton-Clustering Spatial-temporal self-attention with graph convolution Joint-level information Deep clustering SAA-Graph
Luo et al. [115] 2021 Skeleton-Agent ST-GCN for future skeleton prediction Graph connectivity Prediction ST-GCN
Suzuki et al. [116] 2021 Skeleton-Agent Autoencoder for children’s gross motor skills detection Movement patterns Reconstruction Autoencoder
Fan et al. [117] 2022 Skeleton-Agent CycleGAN for enhanced feature extraction accuracy Enhanced features Reconstruction CycleGAN
Jiang et al. [118] 2022 Skeleton-Agent Deep NN for pedestrian behavior detection Keypoint trajectories Trajectory tracking Deep NN
Li et al. [65] 2022 Skeleton-Autoregressive Self-Trained SGCN with iForest initialization Skeleton features Self-trained regression SGCN
Tani & Shibata [119] 2022 Skeleton-Distribution Frame-wise AGCN with Gaussian distribution modeling Action-ness features Distribution modeling AGCN
Yang et al. [107] 2022 Skeleton-Clustering Dual-branch model for humans and objects Skeleton+object features Clustering Dual-branch
Chen et al. [120] 2023 Skeleton-Clustering Multiscale attention-based adjacency matrices Hierarchical graph features Deep clustering MSTA-GCN
Javed et al. [121] 2023 Skeleton-Clustering Decoder-free direct clustering of skeleton features Graph convolution features Direct clustering GCN
Sato et al. [122] 2023 Skeleton-Distribution User cue-guided zero-shot learning framework User cue embeddings Distribution modeling Zero-shot framework
Song et al. [123] 2023 Skeleton-Agent GAN model for global/local motion features Motion features Adversarial learning GAN
Su et al. [87] 2023 Optical Flow Prime framework with optical flow and skeleton fusion Optical flow + skeleton Multi-modal fusion NMS strategy
Yan et al. [124] 2023 Skeleton-Clustering Deep memory clustering with real-time updates Memory features Memory clustering GC-Autoencoder
@@
Newer techniques, such as multi-task learning [125] and causality learning [126, 127, 128, 129], have been explored in RGB-based VAD models, allowing for more specialized categories. This progress is expected to guide future skeleton-based VAD research. In contrast, optical flow removes scene appearance information and is less information-dense than skeletal data. Therefore, it is often used as complementary semantics to RGB sequences and is not typically applied as a standalone VAD approach. Studies have shown that such intermediate modalities, which focus on temporal information, not only mitigate privacy concerns but also avoid enrollment sensitivities similar to those in RGB-based methods. In real-world scenarios, factors such as scene changes, camera angles, and weather conditions cause significant pixel-level variations in color images, leading to high false positive rates in unsupervised VAD methods that rely solely on normal video modeling. Scene-robust VAD has thus become a long-standing challenge. The Desensitized Intermediate Modalities (DIM) approach, which focuses on motion information while ignoring background scenes, allows VAD to concentrate on learning scene-robust video normality. The hierarchical taxonomy of P2VAD methods with DIM is shown in Fig. 4, while the technical comparisons of each method are summarized in Table IV.
Figure 4: Hierarchical taxonomy and related works of P2VAD with Desensitized Intermediate Modalitie (DIM).
@@ V Edge-Cloud Intelligence Empowered P2VAD
TABLE V: Technical Comparison of ECI empowered P2VAD Methods
Method Year Category Core Contributions Privacy Technique System Architecture Key Features
Liu et al. [94] 2020 Privacy Computing System-level data security and privacy preservation Multi-Party Secure Computing Distributed system MPC collaboration
Cheng et al. [91] 2020 Privacy Computing SecureAD framework with additive secret-sharing Homomorphic encryption Deep neural networks Secret-sharing protocols
Giorgi et al. [92] 2022 Edge-Cloud Differential privacy in end-edge-cloud data exchange Differential privacy Edge-cloud architecture Private feature vectors
Wen et al. [136] 2023 Federated VAD Comprehensive survey of federated learning approaches Distributed learning Federated architecture Survey paper
Doshi et al. [95] 2023 Federated VAD FLVAD: Transformer-based federated anomaly detection Federated learning Transformer architecture Local+global models
Al-Dujaili et al. [96] 2024 Federated VAD CLAP: Collaborative Learning with Privacy Federated learning Collaborative framework Pseudo-labeling
Chen et al. [93] 2024 Privacy Computing Federated learning with differential privacy integration Differential privacy Federated system Noise-based protection
Liu et al. [1] 2024 Edge-Cloud Networking systems framework for VAD Secure transmission Edge-cloud integration Hardware+algorithm
Wang et al. [137] 2024 Edge-Cloud End-to-end distributed architecture Edge reasoning End-edge-cloud Task distribution
@@
With the growing adoption of video IoT systems in smart cities and the expansion of online video platforms, data transmission and processing in VAD systems have transitioned from centralized local devices to distributed hybrid architectures that combine edge and cloud computing. Edge-cloud intelligence-enabled P2VAD research focuses on leveraging privacy computing and distributed privacy-preserving techniques to enhance the security of VAD systems across end devices, edge nodes, and the cloud. Unlike the approaches presented in Sections III and IV, which focus on privacy risks during data acquisition and preprocessing by removing identifiable appearance information, Edge-Cloud Intelligence (ECI) empowered VAD emphasizes system-level data security and privacy preservation [94]. Although localized VAD models may still use RGB sequences as inputs, these computational processes typically occur on user-trusted devices, such as smartphones or PCs. Instead of collecting raw data, data centers and cloud servers process encrypted information or intermediate features, ensuring that even in the event of a system breach, sensitive information remains secure. Furthermore, the distributed processing of surveillance videos from different scenes prevents distribution bias caused by simple data aggregation, allowing cloud models to perform cross-scene anomaly detection.
@@
VAD systems for large-scale applications handle massive amounts of video data, which typically need to be transmitted and processed across various edge devices, servers, and data centers. Without effective privacy protection mechanisms, the exchange and integration of information in edge-cloud systems pose significant risks of data leakage and misuse. Therefore, system-level privacy protection not only safeguards user data but also strengthens societal trust in such systems. This chapter explores three major directions: (1) Privacy Computing-enhanced VAD, which reduces the risk of data leakage through privacy computing techniques; (2) Edge-Cloud Collaboration VAD, which focuses on data security and performance optimization during collaboration between end devices and the cloud; and (3) Federated VAD, which explores privacy-secure VAD model training via distributed learning without sharing raw data. These approaches complement one another and collectively improve system-level privacy security. The taxonomy and summary of NIE empowered P2VAD methods are provided in Fig. 6 and Table V, respectively.
Figure 6: Hierarchical taxonomy and related works of End-Cloud Intelligence (ECI) Empowered P2VAD.
@@ V-B Edge-Cloud Collaboration VAD
Edge-Cloud Collaboration VAD is a distributed architecture that strategically divides computing tasks between edge devices and the cloud, aiming to maximize the use of computing resources and alleviate users’ privacy and security concerns about data centers [137]. In this framework, video data captured from real-world scenes can be initially processed on edge devices close to the data source, and only essential features or privacy-protected data are sent to the cloud for further analysis. Since only user-trusted local devices access raw data containing identifiable information, the end-edge-cloud data exchange process and global model training can avoid the risk of data theft and misuse. Existing research in this area is typically categorized into three dimensions: (1) Edge reasoning: smartphones and PCs deployed on the end-side utilize local computing power to preprocess video streams and detect simple anomalies in real-time. Lightweight deep VAD models are often used, reducing data transmission and minimizing privacy exposure; (2) Cloud aggregation: the cloud collects intermediate outputs from the edge, such as hidden features or decision vectors devoid of identifiable information, to train cross-scene models for detecting complex, long-duration anomalies; and (3) Secure data transmission: data is encrypted during transmission across devices, ensuring that sensitive information is not leaked.
@@ VI-A1 Generalized VAD Datasets
TABLE VI: Statistical information of the generalized VAD datasets.
@@
The generalized datasets contain RGB video sequences captured from real-world scenes and are commonly used for traditional deep VAD model training and testing. The details are presented in Table VI. In P2VAD models, these datasets are often encoded or transformed, with morphological abstraction, optical flow extraction, or skeleton detection applied.
@@ VI-B3 Privacy-Preservation Oriented Metrics
  • Degree of De-identification: This measures the effectiveness of techniques like visual desensitization and feature abstraction in removing identifiable information. This can be indirectly assessed through image quality metrics such as Structural Similarity Index Measure (SSIM) or Peak Signal-to-Noise Ratio (PSNR). A higher degree of de-identification corresponds to lower similarity between the processed and original content.
@@ VII Discussion
TABLE VII: Comparison of P2VAD Pathways.
@@
P2VAD with Non-Identifiable Elements In the post-data acquisition step, encryption compression, appearance abstraction, or NVLCs are used to remove personally identifiable information from the RGB sequence. Eliminates privacy concerns and enhances transparency at data source. Encryption compression and appearance abstraction can work on existing videos. Encrypted video reduces redundant background and increases inference speed by over 50 times. NVLCs can function in extreme conditions, such as darkness and high-speed environments. Encryption retains only high-frequency components, limiting detection performance of appearance-motion joint anomalies. Performance of appearance abstraction depends on additional methods (e.g., object detection and instance segmentation). Non-visible light videos require specialized sensing equipment, increasing hardware and data preparation costs. The ideas of RGB-based VAD methods are often incompatible with such data. Efficient video compression can enable the real-time P2VAD system. Lightweight object detection and zero-shot segmentation models (e.g., SAM) can handle complex scenes. Sensing and representation technologies with event cameras and infrared will further P2VAD development. AI technologies (e.g., adversarial samples) can enhance data security and the robustness of P2VAD models.
@@
Sections III-V have presented the implementation and development of various categories of P2VAD approaches that are motivated by different research objectives and underlying assumptions. While existing works have made notable progress in data collection, model learning, and system deployment, significantly advancing privacy-preserving techniques in video anomaly detection, P2VAD, as an emerging interdisciplinary research topic, still faces unresolved challenges in real-world applications. Furthermore, the proliferation of intelligent video surveillance systems and the widespread popularity of short video applications (e.g., TikTok, Kuaishou, and Youtube) among Generation Z provide a broad market for P2VAD. Meanwhile, advancements in AI security and IoT technologies are expected to deeply integrate with VAD research, significantly driving the development of P2VAD. The brief comparison of the strengths, weaknesses, and future directions of various P2VAD Pathways are outlined in Table VII. Moreover, from the perspective of privacy preservation and the research of trustworthy P2VAD systems for large-scale video IoT, this section discusses the challenges and opportunities of P2VAD, analyzing the limitations of existing approaches and trends for future exploration.
@@ VII-A2 Computational Overhead in Heterogeneous Systems
Urban-scale VAD systems often rely on edge intelligence techniques (e.g., mobile computing, edge-cloud collaboration) to optimize the allocation of computational resources and improve overall performance. However, addressing privacy preservation in such large-scale heterogeneous systems—comprising numerous devices such as surveillance cameras, smartphones, local gateways, data centers, and servers—not only requires efficient privacy-preserving computing strategies but also faces significant increases in computational cost. Edge computing typically requires sensitive raw RGB sequences to be processed locally on resource-constrained yet trusted client devices. Additionally, intermediate features often need to be encrypted before transmission to servers or global models to prevent data misuse during transmission. This increases the computational burden on local devices and affects the system’s real-time detection performance. Deploying privacy-compliant VAD systems in scenarios requiring fast response and large-scale application remains a bottleneck, necessitating the development of specialized privacy-preserving computing techniques for video data and efficient offloading algorithms for video IoT devices.
@@ VII-A4 Common Challenges from Emerging AI Technologies
Generative models [152] like large language models (e.g., Generative Pre-Training [153] and Contrastive Language-Image Pretraining [154]) have been introduced to enhance VAD’s ability to understand and describe anomalies [155, 156, 157, 158]. However, their human-like interaction capabilities [159] have further exacerbated public concerns about privacy in VAD. For example, when capturing supplementary semantics in RGB video sequences to improve VAD performance, large models tend to fully describe all visual information, including facial details, clothing appearance, gender, and ethnicity, without considering privacy implications [160]. While spatial appearance and temporal motion cues are essential for VAD tasks, the presence of sensitive identity information is typically irrelevant and exacerbates public distrust. Furthermore, Embodied Artificial Intelligence (EAI) [161] and Explainable Machine Learning (XML) introduce new scenarios and research perspectives for P2VAD. EAI combines robots and perception devices capable of real-time environmental sensing and autonomous responses to emergencies, yet these systems must process large amounts of sensor data (e.g., RGB cameras, infrared cameras, and event cameras), making privacy concerns prominent. XML improves the transparency of P2VAD systems by allowing users to understand the model’s detection process and decision logic. However, it may expose sensitive information, raising questions about balancing explainability and privacy preservation.